OpenStudy (anonymous):

draw the ellipse and find the center, eccentricity, and foci of x^2+4y^2-6x-24y+41=0

OpenStudy (anonymous):

here is a nice picture mertsj will give an explanation, which i imagine involves completing the square to write in standard form http://www.wolframalpha.com/input/?i=ellipse++x^2%2B4y^2-6x-24y%2B41%3D0

OpenStudy (mertsj):

\[x^2-6x+9+4(y^2+6y+9)=-41+45\] \[\frac{(x-3)^2}{4}+\frac{(y+3)^2}{1}=1\] Center (3,-3) a=2, b=1, c=sqrt3 e=sqrt3/2

OpenStudy (anonymous):

useful tool that "completing the square"

OpenStudy (mertsj):

yep

OpenStudy (mertsj):

But I see I got a sign wrong.. It's negative 24 y

OpenStudy (mertsj):

So center is actually (3,3)

OpenStudy (mertsj):

Wish there was some way to correct a post.

OpenStudy (anonymous):

how about the eccentricity?

OpenStudy (anonymous):

the foci,rather,.,.not the eccentricity,.,.

OpenStudy (mertsj):

Foci are positive and negative c units to the right and left of the center. c=sqrt (a^2-b^2)